A portfolio comprises 100 shares of Quicksand Mining Co. ($130 per share) and 200 shares of Bad Bet Bond Fund ($32.50 per share). The expected return on Quicks & Mining Co.’s shares is 60% and the expected return on Bad Bet Bond Fund’s shares is 30%, what is the expected return for the portfolio?
Answer #1

Total value of Quicksand Mining Co. = 100 * 130 = 13,000

Total value of Bad Bet Bond Fund = 200 * 32.5 = 6,500

Total value =13,000 + 6,500 = 19,500

Expected return for the portfolio = (13,000 / 19,500)*0.6 + (6,500 / 19,500)*0.3

Expected return for the portfolio = 0.4 + 0.1

Expected return for the portfolio = 0.5 or 50.00%